OpenStudy (anonymous):

At what rate is a car traveling, if it goes 157.5 miles in 2.5 hours?

OpenStudy (anonymous):

Distance = Rate * Time In this case: Distance = 157.5 Time = 2.5 hours Substituting for the formula: 157.5 = 2.5r Divide both sides by 2.5 to find the rate: r = 63

OpenStudy (anonymous):

i have seen this before it is B

OpenStudy (anonymous):

Thank you @HWBUSTER00

OpenStudy (dmndlife24):

\[rate = \frac{ distance }{ time }\] \[r = \frac{ 157.5 }{ 2.5 }\] \[r = 157.5/2.5 = 63\] \[r = 63\] \[So~the~car~is~traveling~at~a~rate~or~speed~of~63~mph\]

OpenStudy (anonymous):

Thank you @dmndlife24

OpenStudy (anonymous):

your welcome